# Constraints on the Variation of the QCD Interaction Scale $\Lambda_{\text{QCD}}$

### v3

@@ I.1 Sensitivity of the comparison of two optical clock transition frequencies to $X_{e}$
Following Ref. [9], we aim to propose a similar dependence on $X_{e}$ for the ratio of two optical frequencies, by probing the contributions of nuclear mass $m_{N}$ and nuclear radius $r_{N}$ to atomic transition frequencies (these contributions produce mass and field isotope shifts). These contributions are small when compared to the optical transition frequency; however, this mechanism indicates a sensitivity to the parameter $X_{e}=m_{e}/\Lambda_{\text{QCD}}$ from measurements comparing the ratio of two optical transition frequencies. In general, comparing two transition frequencies $\nu_{a}$ and $\nu_{b}$, we obtain
@@
where $K_{\text{MS}}$ and $K_{\text{FS}}$ are the mass and field shift coefficients respectively and $m_{N}$ is the mass of an atom with atomic mass number $A$ (which is largely determined by the nuclear mass $m_{N}$). The mass shift term dominates for light nuclei whilst the field shift term dominates for heavy nuclei. Results have been obtained for the ratio of the optical transition frequencies of $^{1}S_{0}\rightarrow\ ^{3}P_{0}$ transitions in Al+ and $^{2}S_{1/2}\rightarrow\ ^{2}D_{5/2}$ transitions in Hg+ [38, 39, 40] and the ratio of the $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}F_{7/2}(F=3)$ electric octupole (E3) and $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}D_{3/2}(F=2)$ electric quadrupole (E2) transition frequencies in Yb+ [9]. We have previously used Equation (8) along with these results to place constraints on the variation of hadronic parameters, see Refs. [41, 40].
When considering low-mass nuclei, the sensitivity to the parameter $X_{e}$ may be yielded from the dependence of optical transition frequencies on the reduced mass (the normal mass shift)
@@
where the nuclear mass $m_{N}$ is proportional to the nucleon mass, $m_{N}\approx Am_{n}$. The sensitivity to other hadronic parameters comes about from the energy shift due to a finite nuclear size.
These effects are extracted from the measurement of the variation of the ratio of two atomic transition frequencies which have different dependencies on varying parameters. In this case, one may say that the dimensionless parameters appear after dividing the variations of dimensionful parameters by the appropriate atomic units of energy $E_{h}=\alpha^{2}m_{e}c^{2}$, length $a_{B}=\hbar/(mc\alpha)$, etc.
In high-mass nuclei, the relative change in energy (or electronic transition frequency $\nu$) in atoms is proportional to
@@
Setting $c=\hbar=1$, and noting the fact that the nuclear radii in all nuclei are reasonably described by a universal formula $r_{N}=A^{1/3}r_{0}$, where $r_{0}\approx 1.2\ \text{fm}$ is the internucleon distance, we yield
@@
The distance between the nucleons $r_{0}$ is determined by the minimum of the potential in the nucleon-nucleon interaction. In the Walecka model [42] this potential is produced by a vector meson exchange repulsion at short distances and sigma meson exchange attraction at larger distances. Both the vector meson and sigma meson masses are proportional to $\Lambda_{\text{QCD}}$, so the range of these potentials is inversely proportional to $\Lambda_{\text{QCD}}$. Thus $r_{0}\propto\frac{1}{\Lambda_{\text{QCD}}}$ and
$\left(\frac{r_{N}}{a_{B}}\right)^{2}$ $\propto\left(\frac{A^{1/3}m_{e}\alpha}{\Lambda_{\text{QCD}}}\right)^{2}$ (16)
Assuming independent variations, the constraints on the variation of the nuclear charge radius $r_{N}$ presented in these references may be repurposed into constraints on the parameter $X_{e}$, noting that the variation of the $r_{N}^{2}$ is proportional to (see Equation (16))
$\frac{\delta(r_{N}/a_{B})^{2}}{(r_{N}/a_{B})^{2}}\sim 2\left(\frac{\delta m_{e}}{m_{e}}+\frac{\delta\alpha}{\alpha}-\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}\right)\sim 2\left(\frac{\delta X_{e}}{X_{e}}+\frac{\delta\alpha}{\alpha}\right)\,.$ (17)
The numerical constraints obtained from the systems of interest are presented in Table 2.
@@
Al+ Hg+ $(-5.3\pm 7.9)\times 10^{-17}$ [8] $(5.5\pm 8.5)\times 10^{-15}$
Yb+(E3) Yb+(E2) $(-1.2\pm 1.8)\times 10^{-18}$ [10] $(0.25\pm 0.38)\times 10^{-15}$
@@
II Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
Refs. [43, 26] have calculated the sensitivity of the binding energy and nuclear radius in various light nuclei to the different hadron masses. These calculations were extrapolated to heavier nuclei. In such nuclei, the errors produced by such extrapolation were found to be smaller than the errors of direct calculations in heavy nuclei. Assuming that $\Lambda_{\text{QCD}}$ does not vary, these calculations were used to determine that the effects of the variation of the quark mass $m_{q}$ are enhanced roughly $10^{5}$ times in the $8.4$ eV “nuclear clock” transition between the ground and first excited states in 229Th and about $10^{8}$ times in the relative shift of the 0.1 eV compound resonance in 150Sm.
Let us now investigate the implications of interpreting their results as being due to a variation of $\Lambda_{\text{QCD}}$. Using Table III of Ref. [26], the variation of the binding energy $E_{S}$ is related to the variation of the hadron masses by
@@ II Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
which were roughly calculated to be $\Delta\mathcal{E}(m_{N})\approx 12$ and $\Delta\mathcal{E}(m_{\pi})\approx-6$.
The variation of the nucleon mass $m_{N}$ can be related to the variation of the fundamental constants by [14, 44, 45, 46, 47, 48]
[…]
where $m_{q,-}=(m_{u}-m_{d})$ is the mass difference. The dependence of the variation of the nucleon mass on this parameter is weak, and we may neglect it. We will also assume that $\frac{\delta m_{s}}{m_{s}}=\frac{\delta m_{q}}{m_{q}}$, where $m_{s}$ is the mass of the strange quark (this assumption may be motivated by the Higgs mechanism of mass generation). Noting the dependence of the pion mass on $\Lambda_{\text{QCD}}$ and $m_{q}$ from the Gell-Mann-Oakes-Renner relation, $m_{\pi}^{2}\sim\Lambda_{\text{QCD}}m_{q}$ [32], we yield the following result for the variation of the binding energy $E_{S}$
@@
It may be instructive to estimate separately kinetic energy contribution $E_{K}\propto 1/m_{n}r_{N}^{2}$. Using Equation (20), together with $r_{N}=A^{1/3}r_{0}$ and the scaling $r_{0}\propto\Lambda_{\text{QCD}}^{-1}$, the fractional variation is
[…]
implying that in the kinetic energy the $\Lambda_{\text{QCD}}$ term strongly dominates over the $m_{q}$ term.
These results have implications for experiments with the “nuclear clock” transition between the ground and first excited states of the 229Th nucleus, first proposed in Ref. [49], where effects of the variation of fundamental constants are strongly enhanced [50]. This transition has been measured by a range of experiments, see e.g. [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. The most precise measurement gives a transition frequency of $\omega=2,020,407,384,335\ (2)$ kHz ( 8.36 eV) for 229Th dopant ions embedded in a calcium fluoride crystal [60].
Using the results presented above, we may write this transition frequency $\omega$ as a sum of nearly canceling components that have different dependencies on the fundamental constants. The low-lying isomeric state in 229Th is theorised to exhibit a different shape and spin configuration to the ground state [61, 62], resulting in a small energy shift due to the residual strong and electromagnetic effects. As such, we may write $\omega$ as a sum of the electromagnetic energy contribution ($E_{C}$) and strong + kinetic energies contribution $E_{S}$:
@@
where $K_{\alpha}$ may be determined by the relation [50]
@@
In Ref. [63] the sensitivity coefficient $K_{\alpha}$ was expressed via the difference in squared nuclear radii $\delta r^{2}$ and the difference of the electric quadrupole moments $\delta Q$ between the ground and isomeric states. Due to the lack of accurate data for electric quadrupole moments, Ref. [64] used the approximation of a constant nuclear density to express $\delta Q$ via a measured value of $\delta r^{2}$ and obtained $K_{\alpha}$ = $0.82\ (0.25)\times 10^{4}$ (with only the experimental error being presented). However, the approximation of a constant nuclear density is not sufficiently accurate in the case where the contributions of $\delta r^{2}$ and $\delta Q$ have opposing signs. Recent, more accurate measurements of $\delta r^{2}$ and $\delta Q$ [65], using the formula expressing $K_{\alpha}$ in terms of these parameters [63], give $K_{\alpha}$ = $0.59\ (0.23)\times 10^{4}$.
We may use this information, along with the results presented above, in order to estimate the nuclear transition’s sensitivity to changes in the QCD scale $\Lambda_{\text{QCD}}$. In similar fashion to the analysis performed in Refs. [50, 26], Equation (23) implies $E_{C}\approx-E_{S}$ and
@@
The Coulomb interaction contribution $E_{C}$ is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (16). Using Eq. (26), along with the sensitivity of $E_{S}$ to $\Lambda_{\text{QCD}}$ presented in (21), we yield
@@
This estimate for the sensitivity of the nuclear clock transition to variations of $\Lambda_{\text{QCD}}$ may have implications for the measurements performed in Ref. [66], which present constraints on the fractional variation of $\Lambda_{\text{QCD}}$ due to an interaction with ultralight dark matter, taking a rough estimate of the sensitivity coefficient as $K_{\text{QCD}}\sim K_{\alpha}\sim 10^{4}-10^{5}$. Eq. (27) allows one to refine such constraints, which are expected to be significantly improved in future measurements.
This effect is supposed to be extracted from the measurement of the variation of the dimensionless ratio of the nuclear transition frequency to an atomic transition frequency, which is proportional to the atomic unit of energy $E_{h}$. For example, in Ref. [60] the ratio of the nuclear transition frequency to an electron transition frequency in the Sr atom has been measured. The relative variation of the Sr transition frequency is not enhanced, so we may neglect its variation.
III Constraint from Big Bang Nucleosynthesis
@@
Ref. [43] (see previous estimates in Refs. [24, 25, 67, 68] and also [69]) computed how the binding energies in light nuclei respond to shifts of the hadron masses and, assuming the QCD scale $\Lambda_{\text{QCD}}$ is fixed, translated the difference between calculated and observed abundances into a fractional change of the average light-quark mass: $\delta m_{q}/m_{q}=K(0.013\pm 0.002)$, where $K\sim 1$ reflects a theoretical uncertainty of about a factor of 2.
@@
The fact that the central value is non-zero reflects the well-known “lithium problem” - the predicted 7Li abundance exceeds the observed abundance by a factor of 3 to 4. It seems that the lithium anomaly survives every Standard model update to date, and the variation of $\Lambda_{\text{QCD}}$ may solve this problem. The most credible fix within the Standard model is gradual Li destruction in old stars, supported by modern diffusion-mixing simulations but not yet universally accepted. Reviews of the primordial lithium problem - its observational status, nuclear physics inputs and proposed solutions - can be found in Refs. [70, 71].
IV Constraint from the Oklo natural nuclear reactor
The Oklo natural nuclear reactor was a naturally occurring fission reactor that operated about 1.8 billion years ago. Its sustained chain reactions produced a neutron spectrum that covered a very low energy resonance in 149Sm. This resonance, with $E_{r}=0.1\ \mathrm{eV}$, is small compared to the depth of the neutron potential (U=50MeVU=-50\ $\mathrm${MeV}), giving the system a high sensitivity to variations of the fundamental constants. Refs. [72, 73, 74, 75, 76] used this high sensitivity to find limits on the variation of $\alpha$, while Refs. [24, 25, 26] estimated limits on variation of the quark mass.

### v4

@@ II.3 Sensitivity of the comparison of two optical clock transition frequencies to hadronic parameters
Following Ref. [Banerjee2023], we aim to propose a similar dependence on $X_{e}$ for the ratio of two optical frequencies, by probing the contributions of nuclear mass $m_{N}$ and nuclear radius $r_{N}$ to atomic transition frequencies (these contributions produce mass and field isotope shifts). These contributions are small when compared to the optical transition frequency; however, this mechanism indicates a sensitivity to the parameter $X_{e}=m_{e}/\Lambda_{\text{QCD}}$ from measurements comparing the ratio of two optical transition frequencies. In general, comparing two transition frequencies $\nu_{a}$ and $\nu_{b}$, we obtain
@@
where $K_{\text{MS}}$ and $K_{\text{FS}}$ are the conventional mass and field shift coefficients respectively and $m_{N}$ is the mass of the nucleus. The mass shift term dominates for light nuclei whilst the field shift term dominates for heavy nuclei.
Before proceeding to specific numbers, we discuss what dimensionless parameters are involved in these variations. These effects are extracted from the measurement of the variation of the ratio of two atomic transition frequencies which have different dependencies on varying parameters. In this case, one may say that the dimensionless parameters appear after dividing the variations of dimensionful parameters by the appropriate atomic units of energy $E_{h}=\alpha^{2}m_{e}c^{2}$, length $a_{B}=\hbar/(mc\alpha)$, etc.
The nucleon mass dependence in light nuclei originates from the dependence of optical transition frequencies on the reduced mass  [^3] This equation accounts for the normal mass shift only, the specific mass shift depends on the transition and is omitted here.
@@
where the nuclear mass $m_{N}$ is proportional to the nucleon mass, $m_{N}\approx Am_{n}$. Therefore, the dimensionless parameter here is $X_{e}=m_{e}/m_{n}\approx m_{e}/\Lambda_{\text{QCD}}$.
The sensitivity to other hadronic parameters comes about from the energy shift due to a finite nuclear size. In high-mass nuclei, the relative change in energy (or electronic transition frequency $\nu$) in atoms is proportional to
@@
Setting $r_{N}=A^{1/3}r_{0}$, we yield
@@
The variation of $r_{N}^{2}$ is proportional to (see Eqns (26) and (14))
$\frac{\delta(r_{N}/a_{B})^{2}}{(r_{N}/a_{B})^{2}}$ $=2\left(\frac{\delta m_{e}}{m_{e}}+\frac{\delta\alpha}{\alpha}-3.5\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.3\frac{\delta m_{q}}{m_{q}}\right)\,.$ (27)
Now we discuss specific numerical results which have been derived from the ratio of the optical transition frequencies of $^{1}S_{0}\rightarrow\ ^{3}P_{0}$ transitions in Al+ and $^{2}S_{1/2}\rightarrow\ ^{2}D_{5/2}$ transitions in Hg+ [Wineland2007, AlHgdrift, Tang2021, DzubaFlambaumMansour2024] and the ratio of the $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}F_{7/2}(F=3)$ electric octupole (E3) and $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}D_{3/2}(F=2)$ electric quadrupole (E2) transition frequencies in Yb+ [Filzinger2023, Banerjee2023]. In our papers Refs. [FlambaumMansour2023, DzubaFlambaumMansour2024] we used these results to obtain constrains on variation of nucleon mass, nuclear radius, and other hadron parameters. The constraints presented in these references and Eq. (27) may be repurposed into constraints on $\Lambda_{\text{QCD}}$. These constraints are presented in Table 2.
@@
Al+ Hg+ $(-5.3\pm 7.9)\times 10^{-17}$ [AlHgdrift] $(3.2\pm 4.9)\times 10^{-15}$
Yb+(E3) Yb+(E2) $(-1.2\pm 1.8)\times 10^{-18}$ [Filzinger2023] $(7.2\pm 11)\times 10^{-17}$
@@
III Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
Refs. [Wiringa2007, Wiringa2009] have calculated the sensitivity of the binding energy and nuclear radius in various light nuclei to the different hadron masses. These calculations were extrapolated to heavier nuclei. In such nuclei, the errors produced by such extrapolation were found to be smaller than the errors of direct calculations in heavy nuclei. Assuming that $\Lambda_{\text{QCD}}$ does not vary, these calculations were used to determine that the effects of the variation of the quark mass $m_{q}$ are enhanced roughly $10^{5}$ times in the $8.4$ eV “nuclear clock” transition between the ground and first excited states in 229Th and about $10^{8}$ times in the relative shift of the 0.1 eV compound resonance in 150Sm.
Let us now investigate the implications of interpreting their results as being due to a variation of $\Lambda_{\text{QCD}}$. Using Table III of Ref. [Wiringa2009], the variation of the binding energy $E_{S}$ is related to the variation of the hadron masses by  [^4] The contribution of the electromagnetic interaction will be considered below.
@@ III Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
which were roughly calculated to be $\Delta\mathcal{E}(m_{N})\approx 12$ and $\Delta\mathcal{E}(m_{\pi})\approx-6$. Using Eqs. (12) and (13), we yield the following result for the variation of the binding energy $E_{S}$
@@
These results have implications for experiments with the “nuclear clock” transition between the ground and first excited states of the 229Th nucleus, first proposed in Ref. [Peik2003], where effects of the variation of fundamental constants are strongly enhanced [FlambaumTh2006]. This transition has been measured by a range of experiments, see e.g. [Seiferle2019, Masuda2019, Yamaguchi2019, Sikorski2020, Katori2024, ZhangBao2024, Tiedau2024, Elwell2024, Hiraki2024, Zhang2024]. The most precise measurement gives a transition frequency of $\omega=2,020,407,384,335\ (2)$ kHz ( 8.36 eV) for 229Th dopant ions embedded in a calcium fluoride crystal [Zhang2024].
Using the results presented above, we may write this transition frequency $\omega$ as a sum of nearly canceling components that have different dependencies on the fundamental constants. The low-lying isomeric state in 229Th is theorised to exhibit a different shape and spin configuration to the ground state [Litvinova, Minkov2021], resulting in a small energy shift due to the residual strong and electromagnetic effects. As such, we may write $\omega$ as a sum of the electromagnetic energy contribution ($E_{C}$) and strong + kinetic energies contribution $E_{S}$:
@@
where $K_{\alpha}$ may be determined by the relation [FlambaumTh2006]
@@
In Ref. [PhysRevLett.102.210801] the sensitivity coefficient $K_{\alpha}$ was expressed via the difference in squared nuclear radii $\delta r^{2}$ and the difference of the electric quadrupole moments $\delta Q$ between the ground and isomeric states.  [^5] Due to the lack of accurate data for electric quadrupole moments, Ref. [Fadeev2020] used the approximation of a constant nuclear density to express $\delta Q$ via a measured value of $\delta r^{2}$ and obtained $K_{\alpha}$ = $0.82\ (0.25)\times 10^{4}$, with only the experimental error being presented. However, the approximation of a constant nuclear density is not sufficiently accurate in the case where the contributions of $\delta r^{2}$ and $\delta Q$ have opposing signs. Recent, measurements of $\delta r^{2}$ and $\delta Q$ [SafronovaYe], using the formula expressing $K_{\alpha}$ in terms of these parameters [PhysRevLett.102.210801], give $K_{\alpha}$ = $0.59\ (0.23)\times 10^{4}$  [^6] Ref. [SafronovaYe] also investigated the sensitivity of $K_{\alpha}$ to a possible octupole deformation. .
We may use this information, along with the results presented above, in order to estimate the nuclear transition’s sensitivity to changes in the QCD scale $\Lambda_{\text{QCD}}$. In similar fashion to the analysis performed in Refs. [FlambaumTh2006, Wiringa2009], Equation (31) implies $E_{C}\approx-E_{S}$ and
@@
The Coulomb interaction contribution $E_{C}$ is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (27).
Using Eq. (34), along with the sensitivity of $E_{S}$ to $\Lambda_{\text{QCD}}$ presented in (30), we yield
@@
This estimate for the sensitivity of the nuclear clock transition to variations of $\Lambda_{\text{QCD}}$ may have implications for the measurements performed in Ref. [Fuchs2024PRX], which present constraints on the fractional variation of $\Lambda_{\text{QCD}}$ due to an interaction with ultralight dark matter, taking a rough estimate of the sensitivity coefficient as $K_{\text{QCD}}\sim K_{\alpha}\sim 10^{4}-10^{5}$. Eq. (35) allows one to refine such constraints, which are expected to be significantly improved in future measurements.
This effect is supposed to be extracted from the measurement of the variation of the dimensionless ratio of the nuclear transition frequency to an atomic transition frequency, which is proportional to the atomic unit of energy $E_{h}$. For example, in Ref. [Zhang2024] the ratio of the nuclear transition frequency to an electron transition frequency in the Sr atom has been measured. The relative variation of the Sr transition frequency is not enhanced, so we may neglect its variation.
IV Constraint from Big Bang Nucleosynthesis
@@
Ref. [Wiringa2007] (see previous estimates in Refs. [Shuryak2002, FlambaumShuryak, Dmitriev2003, Dmitriev2004] and also [BERENGUT2010114]) computed how the binding energies in light nuclei respond to shifts of the hadron masses and, assuming the QCD scale $\Lambda_{\text{QCD}}$ is fixed, translated the difference between calculated and observed abundances into a fractional change of the average light-quark mass: $\delta m_{q}/m_{q}=K(0.013\pm 0.002)$, where $K\sim 1$ reflects a theoretical uncertainty of about a factor of 2.
@@
The fact that the central value is non-zero reflects the well-known “lithium problem” - the predicted 7Li abundance exceeds the observed abundance by a factor of 3 to 4. It seems that the lithium anomaly survives every Standard model update to date, and the variation of $\Lambda_{\text{QCD}}$ may solve this problem. The most credible fix within the Standard model is gradual Li destruction in old stars, supported by modern diffusion-mixing simulations but not yet universally accepted. Reviews of the primordial lithium problem - its observational status, nuclear physics inputs and proposed solutions - can be found in Refs. [Fields2011, Cyburt2016].
V Constraint from the Oklo natural nuclear reactor
The Oklo natural nuclear reactor was a naturally occurring fission reactor that operated about 1.8 billion years ago. Its sustained chain reactions produced a neutron spectrum that covered a very low energy resonance in 149Sm. This resonance, with $E_{r}=0.1\ \mathrm{eV}$, is small compared to the depth of the neutron potential (U=50MeVU=-50\ $\mathrm${MeV}), giving the system a high sensitivity to variations of the fundamental constants. Refs. [Shlyakhter1976, DAMOUR1996, FUJII2000, Gould2006, Petrov2006] used this high sensitivity to find limits on the variation of $\alpha$, while Refs. [Shuryak2002, FlambaumShuryak, Wiringa2009] estimated limits on variation of the quark mass.